OpenStudy (anonymous):

The dimensions of a rectangular prism are the same as a Pythagorean triangle with legs of 12 and 16. The volume of the prism is

OpenStudy (anonymous):

Hint:USe pythagorean theorem

mathslover (mathslover):

good hint... try to calculate the third side first mrs km0221

OpenStudy (anonymous):

i get the hint but how

OpenStudy (asnaseer):

you cannot find the the volume without knowing the length of the prism

OpenStudy (anonymous):

 h^2=p^2+b^2

OpenStudy (asnaseer):

you can find the area of its base

OpenStudy (asnaseer):

from the information you are given

mathslover (mathslover):

\[{16^2 - 12^2 = b^2}\] use a^2-b^2 = (a+b)(a-b) identity to solve this

OpenStudy (asnaseer):

but there seems to be an inconsistency in your question - "rectangular" prism would normally have a rectangular base - not a triangular base? or maybe the question is badly worded?

mathslover (mathslover):

oh yes... right anjum mrs km0221 confirm your quest. again please

OpenStudy (anonymous):

yeah what ever is going on i do not ge it . i tried to put it in PT but it makes no sense.

OpenStudy (anonymous):

The dimensions of a rectangular prism are the same as a Pythagorean triangle with legs of 12 and 16. The volume of the prism is